The activation energy (Qv) vacancy in pure Ag is 1.762 x 10-19 J/atom. The atomic weight and density for Ag are 107.870 g/mol and 10.5 g/cm3 respectively. Also given the Avogadro’s number is 6.022 x 1023 atom/mol and Boltzmann’s constant is 1.38 x 10-23 J/atom.K.

1.Calculate the value of N, the total number of atomic sites per cubic meter in Ag.

2.Calculate the equilibrium concentration of vacancies (Nv) per cubic meter in pure Ag at 750oC
Answer #1

1) The equilibrium concentration of vacancies per cubic meter is given as:

N_v = Ne^{\frac{-Q_v}{kT}}

where N= {\frac{N_A\rho_{(Ag)}}{A_{Ag}}}

Given \rho_{Ag} = 10.5 g/cm^3

A_{Ag} = 107.87 g/mol

N_A = 6.022\times 10^{23} atoms/mol, and K = Boltzmann’s constant = 1.38 x 10-23J/atom.K.

Hence N= \frac{6.022\times 10^{23}\times (10.5)}{107.87} = 5.86 \times 10^{22 } cm^{-3}

N= \frac{6.022\times 10^{23}\times (10.5)}{107.87} = 5.86 \times 10^{28 } m^{-3}

2) Equilibrium concentration of vacancies (Nv):

N_v = Ne^{\frac{-Q_v}{kT}}

N_v = (5.86\times 10^{28})e^{\frac{1.762\times 10^{-19}}{(1.38\times 10^{-23})(750+273)}}

N_v = 2.23\times 10^{23}m^{-3}